OpenStudy (anonymous):

Identify the 16th term of a geometric sequence where a1 = 4 and a8 = -8,748

OpenStudy (sidsiddhartha):

first term a1=4 and a8=-8748 if r=common ratio then \[a1*(r ^{7})=-8748\] \[r ^{7}=\frac{ -8748 }{ 4 }=-2187\] r=3 now 16th term \[a1*(r ^{15})=4*(3^{15})\]

OpenStudy (anonymous):

so then I multiply and simplify. giving me the answer? or is there more to simplify?

OpenStudy (sidsiddhartha):

yes multiply and simplify

OpenStudy (anonymous):

so would it be -57,395,628? or is my math off?

OpenStudy (sidsiddhartha):

yup

OpenStudy (anonymous):

thanks!